OpenStudy (anonymous):

True or false: tan(pi/2 - x) = cot x

OpenStudy (perl):

Let me give you a list of 'cofunctions' sin( pi/2 - x ) = cos x cos( pi/2 - x ) = sin x sec( pi/2 - x ) = cosec x cosec( pi/2 - x ) =sec x tan( pi/2 - x ) = cot x cot( pi/2 - x ) = tan x

OpenStudy (perl):

note that cos is short for cosine, that indicates its the cofunction of sine cot is short for cotangent, which is the cofunction of tangent